Consider redox reaction: Hg2+(aq) + Sn2+(aq) --->> Hg(I)+Sn4+(aq)
Determine the cell potential (in Volts) at 21oC with the following ion concentrations:
Hg2+(aq) = 0.2907 M
Sn2+(aq) = 1.1006 M
Sn4+(aq) = 0.4066 M
3 decimal places
R= 8.314 J/mol . K
F= 96485 c/mol e-
" 3 decimal places"
1
Expert's answer
2020-10-19T14:04:57-0400

Hg2+(aq) = 0.2907 M


Sn2+(aq) = 1.1006 M


Sn4+(aq) = 0.4066 M


3 decimal places


R= 8.314 J/mol . K


F= 96485 c/mol e-